A circuit consists of a 6 ohm resistor, a 0.3 farad capacitor, and an AC voltage source supplying V(t) = 120 cos(10 t) volts. 1. Write the differential equation for the charge on the capaciton q+0.55 0.56 q=20 cos(10t) X 2. Write a formula for the solution, assuming q(0) 0 q(f) =
Answer #1

Apply the Kirchhoffs loop law to the RC circuit to obtain the loop equation. The expression for the Kirchhoffs loop equatioThen write the linear differential equation of first order for charge as follows: -iR-9(t) +V(t) = 0 (dal) R+90) =v (1) dg()The solution for the above differential equation is, q(t)(IF)= (20 cos (101))(IF)dt+C 9(0)= c., (5 (20cos (10)([F]dt +C) (0)Assume q(0) = 0 to fine out the constant value C. 90-(021( 1. cos (10)-tosin (10) 2012 9(0)=(0.2) ( 155 )cos(1000) +10sin (10